OpenStudy (blackfireskull):

Solve 6x - 2 = 4x + 13 for x. 15/2 11/2 3/12

Vocaloid (vocaloid):

start by adding 2 to each side

OpenStudy (abhisar):

First you bring all the terms with x together.

OpenStudy (blackfireskull):

so 6x+4x?

OpenStudy (abhisar):

6x -2 = 4x + 13 So we want to bring 6x and 4x at one side together and the rest of terms together at another side. So, it will become 6x-4x = 13 +2

OpenStudy (abhisar):

Getting me?

OpenStudy (blackfireskull):

yes. so would that be 10x=15?

OpenStudy (abhisar):

You added 6x and 4x, look carefully. It's -4x so you need to subtract 4x from 6x

OpenStudy (blackfireskull):

oh .-. sorry. a little tired rn. so 2x=15?

OpenStudy (abhisar):

Yes, in the final step divide both sides with 2

OpenStudy (blackfireskull):

so it would be 15/2 :)

OpenStudy (fibonaccichick666):

Tip: Think of this process as reversing the order of operations. SO instead of PEMDAS do Sad Mep

OpenStudy (abhisar):

Correct!